# JMLR: Joint Medical LLM and Retrieval Training for Enhancing Reasoning and Professional Question Answering Capability

### v2

@@ Abstract
With the explosive growth of medical data and the rapid development of artificial intelligence technology, precision medicine has emerged as a key to enhancing the quality and efficiency of healthcare services. In this context, Large Language Models (LLMs) play an increasingly vital role in medical knowledge acquisition and question-answering systems. To further improve the performance of these systems in the medical domain, we introduce an innovative method that jointly trains an Information Retrieval (IR) system and an LLM during the fine-tuning phase. This approach, which we call Joint Medical LLM and Retrieval Training (JMLR), is designed to overcome the challenges faced by traditional models in handling medical question-answering tasks. By employing a synchronized training mechanism, JMLR reduces the demand for computational resources and enhances the model’s ability to leverage medical knowledge for reasoning and answering questions. Our experimental results demonstrate that JMLR-13B (81.2% on Amboos, 61.3% on MedQA) outperforms models using conventional pre-training and fine-tuning Meditron-70B (76.4% on AMBOSS, 60.3% on MedQA). For models of the same 7B scale, JMLR-7B(68.7% on Amboos, 51.7% on MedQA) significantly outperforms other public models (Meditron-7B: 50.1%, 47.9%), proving its superiority in terms of cost (our training time: 37 hours, traditional method: 144 hours), efficiency, and effectiveness in medical question-answering tasks. Through this work, we provide a new and efficient knowledge enhancement tool for healthcare, demonstrating the great potential of integrating IR and LLM training in precision medical information retrieval and question-answering systems.
@@ 1 Introduction
Informed clinical decision-making is fundamentally rooted in understanding and following the logical diagnosis chain. However, the expertise required to employ this diagnostic logic effectively is not universally accessible, posing a significant challenge in healthcare. Recent developments in Large Language Models (LLMs) offer potential solutions to democratize access to this essential aspect of medical knowledge (Achiam et al., 2023; Touvron et al., 2023; Singhal et al., 2022; Peng et al., 2023; Bommasani et al., 2021; Hoffmann et al., 2022; Raffel et al., 2020; Gao et al., 2020; Computer, 2023; Soldaini et al., 2023; Bubeck et al., 2023; Wei et al., 2022; Wang et al., 2022; Yang et al., 2023). LLMs’ historical development and evaluation predominantly focused on generalized tasks, with models pretrained from large amounts of data sourced from various online platforms. This approach, however, often resulted in a dearth of high-quality, domain-specific evidence, a shortfall particularly evident in the medical field Roziere et al. (2023). A viable strategy to expand medical knowledge is to initiate training with medical-purpose LLMs and then refine this training with targeted domain-specific datasets. This hybrid training approach, blending broad linguistic capabilities with specific medical expertise, has been effectively applied in various medical models (Yunxiang et al., 2023; Zhang et al., 2023; Toma et al., 2023; Chen et al., 2023b).
@@
However, LLMs face the significant challenge of “hallucination”, where models generate plausible but incorrect or unverified information (Ji et al., 2023; Bang et al., 2023). Such errors raise serious concerns in healthcare, where precision and reliability are imperative (Ahmad et al., 2023). Hence, recent works proposed retrieval augmented generation (RAG): they first train a retriever to obtain relevant documents from a large corpus based on the input query and then train an LLM to generate a response based on the input query and the retrieved documents (Lewis et al., 2020; Borgeaud et al., 2022a; Cheng et al., 2023; Xiong et al., 2024). By retrieving validated, up-to-date documents, RAG identifies relevant knowledge and provides contextual grounding for LLMs, thereby alleviating hallucination issues (Shuster et al., 2021; Zhu et al., 2023). Although RAG shows high accuracy in benchmarks for open-domain question answering, its effectiveness in specific domains is yet in doubt because retrievers trained in the general domain usually perform worse than those fine-tuned in the specific domain. Fine-tuning retrievers requires document-query pairs, which may not be readily available for a specific domain. Labeling such medical pairs specifically for this purpose incurs additional time and financial costs. Moreover, RAG trains retriever and LLM separately. When training LLM with a frozen retriever, it gathers documents it deems most relevant, which may not always align with what the LLM requires to produce the correct answer.
@@
To solve these challenges, we introduce Joint Medical LLM and Retrieval Training (JMLR), a novel approach that synchronizes the training of an LLM and retriever. We evaluate whether JMLR helps enhance reasoning and professional question-answering capability. Like RAG, JMLR enhances question answering by fetching relevant documents. This involves including each retrieved document in the initial input question before the LLM’s response attempt, thereby augmenting the input to train the LLM for answer generation. In addition to RAG, JMLR introduces a unique mechanism called LLM-Rank loss to train the retriever. This is achieved by evaluating the improvement in LLM’s performance upon including any candidate retrieved documents. We calculate the log probability of the LLM’s answer with each candidate, adopting the negative of these values as the relevance score for each candidate. When reducing this loss, the retriever is trained to prioritize candidate documents that significantly aid the LLM. As shown in Figure 1, JMLR trains LLM and retriever jointly in the same step and thus improves medical question answering, especially in scenarios requiring nuanced understanding and specific information retrieval.
@@
To validate the effectiveness of JMLR in terms of reducing computational resource requirements and fully utilizing given knowledge, we designed three experimental setups. These experiments evaluated whether direct fine-tuning, instead of traditional pretraining, would improve JMLR’s performance. We utilized external resources, including the MIMIC-IV dataset (Johnson et al., 2020), medical textbooks, and diverse medical guidelines for knowledge expansion. Our research presents a novel approach compared to the conventional pretrain-finetune process (initial pretraining followed by finetuning on a downstream task). Our technique has demonstrated superior results, achieving an impressive 53.8% accuracy on the MedQA dataset and 70.2% on AMBOSS, outperforming the traditional method’s 49.8% on MedQA and 52.8% on AMBOSS. Furthermore, our 13B model surpasses the performance of both the Meditron 70B and ChatGPT models in the USMLE and AMBOSS datasets’ question-answering tasks. Moreover, evaluation by both GPT-4 and domain experts supports the superior performance of JMLR. We summarize our key contributions as follows:
We propose a new method 111The code, along with selected retrieval data that can be made public, is included in the supplementary material and will be made publicly accessible upon the paper’s acceptance. that integrates retrieval and LLM training together, resulting in JMLR surpassing the state-of-the-art open source model by 12%.
@@
Our training method decreases the training time required for knowledge expansion to just 37 hours. This is a considerable reduction compared to the conventional pretrain-finetune process, where pretraining alone takes 127 hours and finetuning adds an additional 17 hours.
JMLR, compared to the public SOTA models, is capable of providing enhanced reasoning for improved explainability, and the performance is better than ChatGPT(winning rate is 0.63).

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated a remarkable potential in medical knowledge acquisition and question-answering. However, LLMs can potentially hallucinate and yield factually incorrect outcomes, even with domain-specific pretraining. Previously, retrieval augmented generation (RAG) has limited success in addressing hallucinations. Unlike previous methods in RAG where the retrieval model was trained separately from the LLM, we introduce JMLR (for Jointly trains LLM and information Retrieval (IR)) during the fine-tuning phase. The synchronized training mechanism enhances JMLR’s ability to retrieve clinical guidelines and leverage medical knowledge to reason and answer questions and reduces the demand for computational resources. We evaluated JMLR on the important medical question answering application. Our experimental results demonstrate that JMLR-13B (70.5%) outperforms a previous state-of-the-art open-source model using conventional pre-training and fine-tuning Meditron-70B (68.9%) and Llama2-13B with RAG (54.9%) on a medical question-answering dataset. JMLR-13B (148 GPU hours) also trains much faster than Meditron-70B (42630 GPU hours). Through this work, we provide a new and efficient knowledge enhancement tool for healthcare, demonstrating the potential of integrating IR and LLM training for medical question-answering systems 111The code(https://github.com/believewhat/JMLR-Joint-Medical-LLM-and-Retrieval-Training), along with selected retrieval data that can be made public, is included in the supplementary material and will be made publicly accessible with CC-BY 4.0 license upon the paper’s acceptance..
@@ 1 Introduction
Effective clinical decision-making relies on a logical diagnostic chain, which requires specialized knowledge that isn’t widely accessible, presenting a major healthcare challenge. Large Language Models (LLMs) show promise in making this critical medical knowledge more accessible (Peng et al., 2023; Yang et al., 2023). Traditionally, LLMs have been developed for general tasks using data from diverse online sources, leading to a lack of high-quality, domain-specific information, especially in medicine Wornow et al. (2023); Singhal et al. (2022). Previous researches expand medical knowledge by continue training general domain LLMs with domain-specific datasets. (Yunxiang et al., 2023; Zhang et al., 2023; Toma et al., 2023).
@@
However, LLMs face the significant challenge of “hallucination”, where models generate plausible but incorrect or unverified information (Ji et al., 2023; Bang et al., 2023). Such errors raise serious concerns in healthcare, where accuracy is imperative (Ahmad et al., 2023). Hence, recent works proposed retrieval augmented generation (RAG): they first train a retriever to obtain relevant documents from a general domain corpus based on the input query and then train an LLM to generate a response based on the input query and the retrieved documents (Lewis et al., 2020; Borgeaud et al., 2022a; Cheng et al., 2023; Xiong et al., 2024). By retrieving domain specific documents, RAG identifies relevant knowledge and provides contextual grounding for LLMs, thereby alleviating hallucination issues (Shuster et al., 2021; Zhu et al., 2023).
Although RAG shows high accuracy in benchmarks for open-domain question answering, its effectiveness in specific domains is yet in challenge because retrievers trained in the general domain usually perform worse than those fine-tuned in the specific domain (Gao et al., 2023; Zhao et al., 2022; Thakur et al., 2021). Fine-tuning retrievers requires document-query pairs, which may not be readily available for a specific domain. Labeling such medical pairs specifically for this purpose incurs additional time and financial costs. Moreover, RAG trains retriever and LLM separately. When training LLM with a frozen retriever, the retrieval may not be optimized to align with LLM to produce the correct answer (Asai et al., 2024; Yoran et al., 2023; Rubin et al., 2022). In this study, we introduce a novel method to jointly train retrieval and LLM. Although the method is generalizable, we focus on the evaluation on the important medical domain where the accuracy of the generated content is crucial.
@@
Specifically, we introduce Joint Medical LLM and Retrieval Training (JMLR), a novel approach that jointly train LLM and retriever. As shown in Figure 2, JMLR presents a novel approach compared to the traditional pretrain-finetune process. JMLR enhances question answering by fetching relevant domain-specific documents that maximize the potential of LLMs. This involves including each retrieved document in the initial input question before the LLM’s response attempt, thereby augmenting the input to train the LLM for answer generation. For training jointly, JMLR introduces a unique mechanism called LLM-Rank loss to train the retriever. This is achieved by evaluating the improvement in LLM’s performance upon including any candidate retrieved documents. We calculate the log probability of the LLM’s answer with each candidate, adopting the negative of these values as the relevance score for each candidate. When reducing this loss, the retriever is trained to prioritize candidate documents that significantly aid the LLM. Our results show that the joint training improves medical question answering, especially in scenarios requiring nuanced understanding and specific information retrieval.
@@
To validate the effectiveness of JMLR in terms of reducing computational resource requirements and fully utilizing given knowledge, we designed three experimental setups. These experiments evaluated whether direct fine-tuning, instead of traditional pretraining, would improve JMLR’s performance. We utilized external data resources for retrieval, including the MIMIC-IV dataset (Johnson et al., 2020), medical textbooks, and diverse medical documents for knowledge expansion. We experimented with different LLMs, including PMC-Llama, Meditron, ChatGPT with or without MedRAG. Our JMLR-7B has demonstrated superior results, achieving an impressive 62.45% accuracy, outperforming the traditional method’s 53.2% on the medical QA dataset. Furthermore, our 13B model (69.2%) surpasses the performance of both the open-source medical LLM (Meditron 70B: 68.9%) and closed-source general LLM (ChatGPT: 58.5%). Moreover, evaluation by both GPT-4 and domain experts supports the superior performance of JMLR. We summarize our key contributions as follows:
We propose a new method that jointly train retrieval and LLM, resulting in JMLR surpassing the state-of-the-art open source model.
By training jointly, we do not require additional human annotation, thereby enhancing flexibility in domain or downstream task adaptation.
@@
When training 7B parameter model with a single Nvidia A100 GPU, our training method decreases the training time required for knowledge expansion to just 100 hours. This is a considerable reduction compared to Meditron’s pretrain-finetune process, where pretraining alone takes 588 hours, and finetuning adds an additional 36 hours.
2 Problem Formulation & Traditional Method
Given a set of medical questions and a set of medical documents containing medical knowledge, our goal is to construct a language model that can provide accurate answers.
Formally, for each question , there is a corresponding correct answer within a set of options , where . The model maps each question to a predicted answer:
where represents the parameters of the model.
Our objective is to find the optimal parameters that minimize the loss function , which measures the discrepancy between the predicted answer and the correct answer :
The loss function can be instantiated as a cross-entropy loss for classification tasks, where the number of classes is equal to the number of options.
To solve this problem, previous methods pretrained LLM to learn medical knowledge for a medical LLM, and then finetuned medical LLM on medical QA task (Chen et al., 2023b; Wu et al., 2023; Yunxiang et al., 2023; Toma et al., 2023). Specifically, such traditional pretrain-finetune pipeline first continued pretraining general domain LLMs on medical documents with next-token prediction loss function, and then finetuned the medical LLMs to select given question and option description with loss function .
3 Joint Medical LLM and Retrieval Training (JMLR)
In comparison to traditional method, JMLR did not pretrain LLM on medical documents. Instead, JMLR selected question-related medical documents as additional context during fine-tuning of an open-domain LLM. Specifically, given a question , we trained retriever to find helpful medical documents from , which was then used to train LLM to generate the final answer .
3.1 Retriever
3.2 LLM
We adapted Llama as our LLM model. The original Llama model supports only a maximum of 4k tokens, which poses a challenge when multiple documents are retrieved. To address this, we utilized the Shifted Sparse Attention (S2-Attn) mechanism (Chen et al., 2023a), which can be used to process long sequences by mitigating the high memory cost and slow processing time associated with standard self-attention in LLMs. S2-Attn divides input into sequence spans within self-attention modules and introduces shifted patterns for inter-span information exchange. This extends context length efficiently without extra computational costs and achieves near-baseline performance. It maintains the standard self-attention architecture during inference, ensuring compatibility with existing infrastructure and optimizations.
3.3 JMLR Architecture
However, Colbert matches queries and documents based on the retriever similarity, returning scores and documents. To enable the retriever system to return documents that are more helpful for providing accurate answers to Llama, we constructed the JMLR architecture. Let be a set of queries, be a corpus of domain specific documents, and be the set of options. Our task is to construct a function parameterized by , which maps a question and a set of documents to an answer .
The objective is to simultaneously learn the optimal parameters and that minimize the combined loss on the retrieved documents and the generated answers, formalized as: where is a loss function that evaluates the correctness of the answer and the relevance of the retrieved documents. The retrieval function is defined by:
and the similarity function scores how relevant each document is to the question .
The combined optimization involves updating both and through gradient descent to minimize the expected loss:
where denotes the gradient of the loss function with respect to both sets of parameters. This dual-parameter optimization ensures that the retrieval function is aligned with the needs of the answer function, enhancing both the accuracy of the responses and the relevance of the information retrieved. To solve this, JMLR structure employs a rank loss, , which links the Llama and retriever systems. Llama’s content quality influences the updating of retriever parameters. If the returned document reduces Llama’s loss function more than document , it indicates that document is more helpful for answering than document . In calculating the rank loss, we incorporate the scores generated by the retriever system. However, due to the large variance in scores, we normalize them to obtain . The essence of is that the retriever system’s score rankings should be updated based on the ranking. in .
where is calculated as:
and represents the contrastive logits:
with the associated decision-based loss differences:
For JMLR model training, the dataset comprises QA pairs, with the input being a question and the output being the correct choice, including the reasoning part. The top 7 documents with the highest scores are aggregated and fed into the model during the training process. Prior to training, we extracted the top 30 documents based on scores from the retriever. During each iteration, 30 scores are generated by the retriever, and we perform weighted random sampling based on these scores, with higher scores having a greater probability of selection. This approach ensures that each document inputted into the LLM model is dynamic. The LLM will likely encounter different useful documents and some that are not useful. This step enhances the LLM’s robustness, helping it discern which knowledge from which documents is useful and which documents are irrelevant.
4 Experiment
4.1 Dataset
Medical document:
Data quality of medical documents is important for LLMs to answer medical questions. Previous work primarily sourced from a wide array of medical research articles and clinical guidelines. Research papers, like those found in PubMed, provide foundational and current information on healthcare. Clinical guidelines are thoroughly designed protocols developed to assist clinicians in making decisions given patient medical history. Other high-quality medical documents to retrieve include medical textbooks.
| Corpus | #Doc. |
| PubMed | 23.9M |
| Textbooks | 18 |
| Cancer Care Ontario | 87 |
| Center for Disease Control and Prevention | 621 |
| Canadian Medical Association | 431 |
| International Committee of the Red Cross | 49 |
| National Institute for Health and Care Excellence | 1.7k |
| Strategy for Patient-Oriented Research | 217 |
| World Health Organization | 223 |
| WikiDoc | 33k |
Medical QA:
The MedQA dataset Jin et al. (2021), sourced directly from the official USMLE website, included a range of sample questions for Step1, Step2CK, and Step3, released between June 2022 and March 2023. These questions represent the complex medical knowledge and ethical scenarios that medical students and practitioners are expected to navigate.
Additionally, Amboss question bank is a comprehensive resource widely used by medical professionals and students. This dataset provided an extensive array of Step1, Step2CK, and Step3-type questions, further enriching our training material with practical and diverse medical scenarios.
The MedMCQA dataset comprises 194k multiple-choice questions from Indian medical entrance exams, covering 2.4k healthcare topics and 21 medical subjects (Pal et al., 2022).
MMLU-Medical is selected from the MMLU dataset Hendrycks et al. (2020), focused on nine subjects most pertinent to medical and clinical knowledge — high school biology, college biology, college medicine, professional medicine, medical genetics, virology, clinical knowledge, nutrition, and anatomy. Since this dataset has no training set, we opt to evaluate it using the LLM originally trained on MedMCQA.
We individually apply supervised finetuning on each QA dataset and subsequently assess their performance on the corresponding test sets, unless otherwise specified. MedQA and Amboss both offer detailed reasoning traces, such as extensive answers or explanations.
4.2 Baselines
For comparison with traditional domain pretraining method, we select several public available medical domain pretrained LLMs as baselines: (1) PMC-Llama, a Llama model further pretrained on medical research articles from PubMed Central (Wu et al., 2023); (2) Meditron, a LLM pretrained on clinical guidelines and research papers (Chen et al., 2023b).
To validate the benefit of joint training, we also compare JMLR with its naive version: RAG, which freezes the retriever and fetches the same document given a specific question.
4.3 Training Details
Both our training process and the conventional finetuning approach employ the AdamW optimizer, with , , and . We implement a cosine learning rate schedule, incorporating a warmup phase that accounts for 10% of the training duration and decays the learning rate to 10% of its peak value. In alignment with the practices outlined in Llama 2-chat Touvron et al. (2023), our training employs a learning rate of , a weight decay factor of 0.1, and manages a batch size of 2. The finetuning phase spans 5 epochs for all iterations. However, we apply a distinct learning rate for ColBERT, set at . The optimization strategy for training ColBERT mirrors that used for Llama 2. Throughout these experiments, we utilize four A100 80 GB GPUs. The total training duration for our method amounts to 25 hours. Additionally, the pretraining phase on medical guidelines takes approximately 138 hours, while the finetuning phase is approximately 17 hours on 7B Llama. This comprehensive training approach ensures the effective adaptation and optimization of the model for specific medical applications.
5 Main Results
| Model | Parameter | Open | MedQA | Amboss | MMLU | MedMCQA | Average |
| Access | |||||||
| GPT-4 | ? | No | 74.7 | 82.1 | 88.4 | 69.5 | 78.6 |
| ChatGPT | ? | No | 50.2 | 49.1 | 69.4 | 51.0 | 54.9 |
| Meditron | 70B | Yes | 60.7 | 76.4 | 73.6 | 65.1 | 68.9 |
| RAG | 13B | Yes | 59.9 | 76.9 | 69.9 | 64.2 | 67.7 |
| JLMR | 13B | Yes | 62.5 | 81.2 | 72.8 | 65.5 | 70.5 |
As shown in Table 2, our JLMR with 13B parameters outperforms previous SOTA open access model Meditron 70B parameters and even closed access model such as ChatGPT. In the following sections, we will explain with more detailed analysis.
5.1 Domain Retrieval > Domain Pretrain
| Model | MedQA | Amboss | MMLU | MedMCQA | Average |
| Llama 2 | 44.0 | 46.5 | 56.3 | 54.4 | 50.3 |
| PMC-Llama* | 42.4 | 43.7 | 59.7 | 57.6 | 50.8 |
| Meditron* | 47.9 | 50.1 | 55.6 | 59.2 | 53.2 |
| RAG# | 54.6 | 70.7 | 62.1 | 62.4 | 62.5 |
| JLMR# | 56.2 | 68.3 | 64.3 | 62.4 | 62.8 |
Domain retrieval models (both RAG-7B and our JMLR-7B) outperform domain pretrained LLMs across all medical QA benchmarks, as shown in Table 3. In the MedQA and Amboss datasets, where JMLR-7B not only outperforms baselines but does so with a notable margin, scoring 56.2% in MedQA and 71.2% in Amboss. This is particularly significant when compared to Meditron-7B, which scores 47.9% and 50.1% in these datasets, respectively. The trend continues in MMLU-Medical and MedMcQA datasets, JMLR-7B achieved scores of 64.3% and 62.4%, respectively, surpassing the scores of PMC-Llama-7B (59.7% and 57.6%), and Meditron-7B (55.6% and 59.2%). Overall, the results illustrate that JMLR-7B model, on average, achieves about a 14% improvement in performance over its closest competitor, Meditron-7B. This highlights the effectiveness of our training method and the substantial advancements it brings to the field of medical benchmarking.
The enhanced performance of JMLR-7B can be partially attributed to its use of Llama-2 as the foundational model, showcasing significantly higher average performance than other pre-trained baselines. However, the distinct edge comes from our integrated training approach that combines retriever and LLM. This methodology not only further enhances Llama-2’s performance in medical benchmarks but also equips the retriever component with the ability to effectively source relevant and beneficial documents to aid the LLM in answering questions. In contrast, ChatGPT, when devoid of medical guidelines, has been known to make basic mistakes, such as recommending vaccinations to pregnant women—a practice that contradicts medical norms as documented in Table 5. Our model, with the support of guidelines, avoids such errors. This synergistic approach substantially reduces hallucinations and bolsters the model’s overall reliability.
5.2 Joint Training > RAG
| Llama-7B | Llama-7B-ColBERT | FT-Llama-7B | FT-Llama-7B-ColBERT | JMLR-7B-Seperate | JMLR-7B | |
| MedQA | 24.98 | 24.01 | 44.00 | 40.60 | 55.30 | 56.20 |
| Amboss | 17.10 | 35.90 | 46.50 | 45.60 | 71.00 | 71.20 |
During fine-tuning, synchronously updating both the Retriever and LLM yields better results than updating only the LLM. Table 2 shows that JLMR models significantly outperform traditional RAG methods. For example, JLMR-13B’s scores (72.8% on MMLU, 65.5% on MedMcQA) is far superior to RAG-13B (69.9% on MMLU, 64.2% on MedMcQA). This result suggests that without updating the retriever, the documents retrieved may be relevant to the question but not necessarily helpful to the LLM’s response. Joint training ensures that the retriever learns which documents are beneficial for Llama-2.
To evaluate the adaptability of the JLMR retriever, we further applied it to other LLMs. Specifically, we employed different retrievers to identify relevant documents, which were then provided to ChatGPT for generating the final answer. The retriever trained with JMLR attained an average accuracy of 57.5%, while the retriever from RAG achieved an accuracy of 56.7%. These results indicate that the JLMR retriever effectively identifies documents that not only support Llama but also aid ChatGPT in selecting the correct answers. Interestingly, this experimental outcome aligns with recent studies on weak-to-strong generalization Burns et al. (2023), such as the Aligner Ji et al. (2024). It appears that training a retrieval module using a weaker model through the JMLR process enables it to function as a model-agnostic plug-and-play module, which subsequently enhances the performance of a stronger model. This suggests a promising research direction for future studies.
6 Ablation Study
To assess the effects of fine-tuning and information retrieval on model performance, we structured our experiments as follows:
Original Model Evaluation: Test the original Llama 2-7B on the MedQA and Amboss datasets.
Retrieval Augmentation: Evaluate the addition of a retriever during the inference stage to assist Llama 2-7B (Llama-7B+Colbert) in answering questions, analyzing the impact of the Colbert retriever.
Fine-Tuning: Fine-tune Llama 2-7B on the Amboss and MedQA training sets to develop FT-Llama-7B.
Retrieval with Fine-Tuning: Add a retriever during the inference stage to assist the fine-tuned model (FT-Llama-7B+ColBERT), to explore the effects of Colbert on the fine-tuned model.
Asynchronous Training: Fix the parameters of the retriever while only updating those of Llama, then reverse to update only the retriever’s parameters using rank loss, resulting in JMLR-7B-Separate.
Joint Training: Train the retriever and Llama simultaneously (JMLR-7B) to compare against separate training and to examine the impact of joint training on enhancing model performance.
Table 4 displays the varied performance metrics of the models at different experimental stages. Observations are as follows:
Llama-7B+Retriever (24.01% on MedQA, 35.90% on Amboss) underperformed compared to the original Llama 2-7B (24.98% on MedQA, 17.10% on Amboss), suggesting that the original Llama 2-7B model struggles to effectively utilize external documents, with longer documents potentially leading to misunderstandings or erroneous information synthesis.
The FT-Llama-7B (44.0% on MedQA, 46.5% on Amboss) showed significant improvements over Llama-7B, indicating that the fine-tuning phase notably enhanced Llama’s capabilities in medical QA. However, this phase did not effectively leverage medical guidelines.
The JMLR-7B-Separate model achieved superior performance (55.3% on MedQA and 71.0% on Amboss), surpassing the FT-Llama-7B-ColBERT (40.6% on MedQA, 46.5% on Amboss). The distinctive fine-tuning approach of JMLR-7B-Separate, which integrates medical documents during the tuning phase, enables the model to utilize medical resources more effectively.
Finally, JMLR-7B (56.2% on MedQA, 71.2% on Amboss) uses a joint training approach that facilitates the retrieval of highly relevant medical documents, enhancing the model’s ability to accurately answer questions by presenting similar case-based treatment plans for consideration.
As illustrated in Table 5, JMLR retrieves closely related cases, significantly improving the model’s decision-making process by offering comparable treatment options.
7 Discussion: JMLR Effectiveness of the reasoning ability
To verify that our model can accurately answer questions and provide detailed reasoning, we tasked JMLR-13B with generating reasons for its responses to the Amboss and MedQA test sets. Similarly, we had ChatGPT generate corresponding reasons. We employed two methods for scoring these reasons: 1) We used GPT-4 to judge which model’s generated reasoning was better based on the ground truth. 2) We engaged three medical professionals to review and assess a sample of 20 questions drawn from the test sets. These doctors were asked to judge which model’s reasoning was closer to the actual, clinically accurate reasons.
In the assessment using GPT-4, JMLR-13B achieved a win rate of 0.63, while under the evaluation of medical professionals, our model attained a win rate of 0.60. These results suggest that simply increasing the model’s parameters is not the most effective way to expand medical knowledge. Instead, JMLR-13B’s ability to consult different medical guidelines for various questions and to find corresponding answers and explanations significantly reduces hallucinations and greatly enhances the model’s explanatory power. This achievement highlights the model’s proficiency in accessing and applying relevant medical information contextually, thereby providing more accurate and credible explanations. The close alignment of the win rates from both GPT-4 and medical professionals’ evaluations further underscores the model’s robustness and reliability in medical question-answering and reasoning tasks.